# Evidence of social learning across symbolic cultural barriers in sperm whales

### v2

@@ Abstract
We provide quantitative evidence suggesting social learning in sperm whales across socio-cultural boundaries, using acoustic data from the Pacific and Atlantic Oceans. Traditionally, sperm whale populations are categorized into clans based on their vocal repertoire: the rhythmically patterned click sequences (codas) that they use. Among these codas, identity codas function as symbolic markers for each clan, accounting for 35-60% of codas they produce. We introduce a computational method to model whale speech, which encodes rhythmic micro-variations within codas, capturing their vocal style. We find that vocal style-clans closely align with repertoire-clans. However, contrary to vocal repertoire, we show that sympatry increases vocal style similarity between clans for non-identity codas, i.e. most codas, suggesting social learning across cultural boundaries. More broadly, this subcoda structure model offers a framework for comparing communication systems in other species, with potential implications for deeper understanding of vocal and cultural transmission within animal societies.
@@ 1 Introduction
Cultural transmission is defined as the transmission of information or behaviors between individuals of the same species by means of social learning [1]. While humans represent a benchmark of such capacity, cultural transmission has been observed in a wide variety of animals, including cetaceans [2, 3], songbirds [4], non-human primates [5], and insects [6]. It typically takes one of three forms: vertical transmission, from adult kin to young kin; oblique transmission, from unrelated adults to young; or horizontal transmission, from peer to peer [7].
@@
Figure 1: Statistical modeling of subcoda structure in sperm whales. A Sperm whale communication consists of rhythmic sequences of clicks, called codas. A coda is specified by a sequence of inter-click intervals (ICIs). Codas are classified into types based on their rhythmic pattern, which can have various degrees of regularity (e.g., 4R2 vs 1+3). B Social groups of sperm whales employ specific vocal repertoires: the set of coda types they use and their associated usage frequencies. As an illustration, we show those of the EC1 and EC2 clans from the Dominica dataset [12]. Only the most numerous coda types are shown: the rest of the vocal repertoires consists of more coda types with residual frequencies. C The subcoda structure can be modeled by considering rhythmic variations within codas of the same coda type. To do so, codas are represented as sequences of discrete inter-click intervals (dICIs), by discretizing absolute ICIs into discrete bins, which can then be considered akin to symbols (e.g. A, B, C, …), providing a tokenization of codas. Different instances of a single coda type can correspond to slightly different dICI sequences. The resulting dICI sequences are modeled using variable-length Markov chains, which can be represented as subcoda trees. These trees can be built for an individual speaker or for a group of speakers, and capture the statistical and memory structures of rhythm variations within codas and in the transitions between those. In other words, the tree captures a vocal style—how they say what they say. The vocal styles of different groups of sperm whales can be quantitatively compared by calculating a distance between their subcoda trees.
Sperm whales live in multi-tiered societies and have a complex vocal communication system [13]. They communicate through rhythmic patterns (codas) of short broadband sounds (clicks), which have traditionally been classified into a finite set of coda types based on the total number of clicks, their rhythm, and their tempo [14, 15] (Fig. 1A). For example, the 4-regular (4R2) type refers to a pattern of four evenly spaced clicks, whereas the 1+3 type refers to two clicks separated by a longer pause followed by two clicks in quick succession. Coda types are thus standardized rhythmic patterns, but individual vocalizations of a given coda type exhibit micro-variations around that pattern.
@@
By applying our modeling framework to acoustic data from the Atlantic and Pacific Oceans, we obtain two main results. First, we partition sperm whale populations into vocal style-defined clans, which we find to recapitulate the previously defined vocal repertoire clans. This confirms that our method does capture meaningful speech characteristics. Second, and crucially, we find that the vocal style of non-ID codas is more similar for more sympatric clans, i.e. clans whose territory overlaps more spatially. In contrast, we do not find an effect of sympatry on the similarity of vocal styles when studying only ID codas. This suggests that geographic overlap induces vocal styles to become more similar between clans, without jeopardizing each clan’s acoustic identity signals. Our results strengthen previous results on the use of ID codas as symbolic markers, while supporting cultural transmission and social learning of vocalizations among whales of different clans, as predicted by theoretical models [22].
@@
Click
Sharp broadband pulses [23]
Inter-click interval (ICI)
Time interval between two consecutive clicks
Coda
Stereotyped, short series of clicks produced by sperm whales in social situations [14]
Coda type
Categorical classification of codas based on rhythm and tempo of coda inter-click intervals [16]
Coda frequency
Usage frequency of each coda type
Coda sample
Set of codas recorded during one event
Coda usage
Set of coda types used by a group of whales
Identity codas
Subset of coda types most frequently used by a sperm whale clan; canonically used to define vocal clans.
Non-identity codas
Coda types produced by whales belonging to multiple vocal clans (i.e., all coda types that are not identity codas)
Vocal clan
Higher-order level of social organization delineated based on social units that share a substantial part of their vocal repertoire [11, 20]
Vocal repertoire
Coda usage and frequency combined. Description of what sperm whales say.
Vocal style
Encodes the rhythmic variations within codas. Description of how sperm whales emit codas
Vocal identity
Concept comprising both vocal style and vocal repertoire
@@ Results
We model the internal structure of codas, in terms of rhythmic variations at the level of clicks, by using variable length Markov chains (VLMCs). Our analytical pipeline is illustrated in Fig. 1C. We build each VLMC in two main steps. We first convert codas, naturally represented as sequences of continuous, absolute, inter-click intervals (ICIs), to sequences of discrete ICIs (dICIs), by discretizing time into bins. In this way, each dICI represents a narrow range of possible ICI values. The bins have a fixed width (or resolution) $\delta t$ and thus implicitly correspond to the temporal resolution of our representation (see Methods for details on the optimal choice of $\delta t$). Note that although ICIs have units of time (seconds), dICIs are (unit-less) symbols (e.g. A, B, C, etc.), representing multiples of $\delta t$ (and so the smaller $\delta t$, the more the symbols). For example, the shortest ICIs will be mapped to the symbol A whereas longer ones will be mapped to symbols further down the alphabet. Hence, each coda (a sequence of ICIs) is mapped to a sequence of discrete symbols (a sequence of dICIs). The second part of the pipeline focuses on modeling the internal structure of codas in terms of dICI sequences. Essentially, we want to estimate transition probabilities from a dICI sub-sequence to the next dICI (Fig. 1B). A standard way would be to describe this using $k$-order Markov chain models, which encode information on previous sub-sequences up to $k$ steps in the past of the sequence. However, it is possible that different sequences of dICIs contain different amounts of information or memory regarding potential next dICIs. This is akin to what happens with words (e.g., a word beginning with “re” can continue in more ways than one starting with “zy”). To account for this possibility while also retaining only the most compressed statistical representation of how codas are structured in terms of dICIs, we employ VLMCs.
@@
Figure 4: Clan overlap influences non-ID coda vocal styles Comparing the similarities of different VLMC models fit for each Pacific Ocean clan for both ID and non-ID coda samples. The y-axis represents the measured distance between the subcoda trees, and the x-axis shows the geographical clan overlap (as calculated in [11]). Each point represents a pairwise comparison between two clans. The effect of overlap on ID coda vocal style similarity is minimal and non-significant while the opposite is true for non-ID codas: overlapping clans produce non-ID codas with a more similar vocal style. The VLMC distances are also typically much greater for ID codas than for non-ID codas. Note that while these results are visually opposite to those reported in Hersh et al. [10], they support the same final conclusions (see “Identity and non-identity codas show different trends” for details).
@@
We find the exact opposite effect when considering vocal style. The similarity in vocal style for ID codas across clans does not depend on the level of clan overlap (Fig. 4a). In contrast, the similarity in vocal style for non-ID codas displays a clear and significant increase (i.e., decreasing subcoda tree-distance) as clan spatial overlap increases (Fig. 4b). In the Supplementary Materials (see Section $2.4.2$), we show that the same results hold at the single coda type level, in addition to the whole clan level , along with an analysis of the confidence intervals. These results imply that the internal structure of codas is more similar for groups that likely spend more time in the same space, akin to accents aligning in human populations that share the same territory [25, 26]. This also highlights the complementarity of vocal repertoire and style: the trends are different precisely because the two concepts describe different aspects of whale speech.
@@ Discussion
For ID codas, we show that the similarity between clan vocal styles is not affected by spatial overlap, while it has recently been shown that the similarity between clan vocal repertoires decreases with overlap [11]. This means that spatial overlap does not affect how whales produce ID codas (in terms of their fine-scale rhythmic structure; our results) but does affect how often they produce them. In contrast, for non-ID codas, we show that the similarity between vocal styles increases with spatial overlap between two clans, while no change was observed for vocal repertoires in previous work on the same dataset. In other words, increasing spatial overlap is correlated with more similar fine-scale rhythmic structure of non-ID codas produced by whales from different clans (our results), but does not affect how often non-ID codas are produced. Our study thus supports and nuances the results of Hersh et al. [11]. We provide further support for selection acting to produce unambiguous, recognizable identity signals in the ID codas. However, ID codas only account for 35% of the total vocalizations; the remaining 65% of codas have traditionally been lumped into a catch-all category (i.e., non-ID codas) and their function remains enigmatic (these numbers are an average over the Pacific clans, and go up to 93% for non-ID codas when counting number of coda types instead of number of codas emitted, see SM 1.1 for details). We could still discriminate among clans using non-ID coda vocal style; however, the increased similarity of non-ID coda vocal styles between clans with greater spatial overlap, as demonstrated here, suggests that non-ID codas are likely vocal cues and not identity signals like the ID codas. Accordingly, vocal repertoire and vocal style capture different and complementary information on sperm whale communication, and should be considered in tandem in future studies.
@@
The most parsimonious explanation for the observed similarity of non-ID coda vocal styles of clans with increasing spatial overlap is social learning across clan boundaries. This is remarkable, given that sperm whale clans belonging to different clans have rarely been observed physically interacting at sea [3]. However, that does not preclude the possibility that they are within acoustic range of each other [33] and that cross-cultural social learning opportunities arise. This explanation is compatible with (and bolsters) past work suggesting that ID and non-ID codas function differently in sperm whale communication, and further suggests that they experience different evolutionary pressures [22]. Whether social learning has facilitated stochastic (i.e., cultural drift) or deterministic (i.e., cultural selection) processes is more difficult to determine, and it is unclear whether the observed non-ID coda vocal style alignment has been neutral or adaptive [34, 28]. Importantly, these findings suggest that vocal learning in sperm whales may not be limited to vertical transmission from related adults to young kin, but that horizontal and/or oblique social learning from outside the natal social unit might also be occuring.
Vocal identity in sperm whales is thus consistent with both cultural selection on ID codas to maintain discrete signals for vocal recognition in sympatry, and social learning between clans leading to a vocal style more similar to that of other whales with which they are in acoustic contact more frequently. This highlights a more complex system of transmission in which clan identity is maintained through selection, while gradual change over time may occur within and across clans for vocalizations which do not function in social recognition and thus may create similar vocal styles.
@@
Our results can be expanded in multiple ways in future work. The first, and the simplest conceptually, would be to conduct the present analysis on a larger dataset. More codas would improve the quality of the statistical analyses and ensure that all codas are represented in realistic proportions for each clan. Moreover, longitudinal datasets might provide direct evidence to discriminate between the social learning hypothesis and competing ones (e.g. drift in vocal style). Similarly, confirmations could emerge from large scale genetic datasets addressing the issues of phylogenetic relatedness (or lack thereof) in clans that are closer in vocal style distance. Such datasets do not exist at present, but efforts towards automated and semi-automated collection techniques are underway (e.g. Project CETI [35]). Second, from a methodological perspective, we could add spectral information (in terms of acoustic frequencies) to the temporal information currently used. Although sperm whale acoustic communication seems mostly based on rhythm, spectral features of individual clicks may convey additional information. This possibility could be incorporated into our method by labeling the dICIs according to the frequency content of the associated click (or by extending the available “alphabet” for the VLMC). Third, it would be interesting to investigate in more detail the function of non-ID codas. Indeed, even though ID codas were only recently formally named for the first time, they have been the primary focus of sperm whale coda research for decades. As previously mentioned, non-ID codas are a catch-all category for anything that is not an ID coda, but that does not mean that all non-ID codas function in the same way. To start to unveil their function, we need to consider the context (behavioral, environmental, etc.) in which different non-ID codas are produced [36]. The pattern we documented may or may not apply to all non-ID codas, but it is at least strong enough that we detect the relationship with clan spatial overlap when collectively considering all non-ID codas.

### v3

@@ Abstract
We provide quantitative evidence suggesting social learning in sperm whales across socio-cultural boundaries, using acoustic data from the Pacific and Atlantic Oceans. Traditionally, sperm whale populations are categorized into clans based on their vocal repertoire: the rhythmically patterned click sequences (codas) that they use. Among these codas, identity codas function as symbolic markers for each clan, accounting for 35-60% of codas they produce. We introduce a computational method to model whale communication, which encodes rhythmic micro-variations within codas, capturing their vocal style. We find that vocal style-clans closely align with repertoire-clans. However, contrary to vocal repertoire, we show that sympatry increases vocal style similarity between clans for non-identity codas, i.e. most codas, suggesting social learning across cultural boundaries. More broadly, this subcoda structure model offers a framework for comparing communication systems in other species, with potential implications for deeper understanding of vocal and cultural transmission within animal societies.
@@ 1 Introduction
Cultural transmission is broadly defined as the transmission of information or behaviors between individuals of the same species by means of social learning [1, 2]. While humans represent a benchmark of such capacity, cultural transmission has been observed in a wide variety of animals, including cetaceans [3, 4], songbirds [5], non-human primates [6], and insects [7]. It typically takes one of three forms: vertical transmission, from adult kin to young kin; oblique transmission, from unrelated adults to young; or horizontal transmission, from peer to peer [8].
@@
Figure 1: Statistical modeling of subcoda structure in sperm whales. A Sperm whale communication consists of rhythmic sequences of clicks, called codas. A coda is specified by a sequence of inter-click intervals (ICIs). Codas are classified into types based on their rhythmic pattern, which can have various degrees of regularity (e.g., 4R2 vs 1+3). B Social groups of sperm whales employ specific vocal repertoires: the set of coda types they use and their associated usage frequencies. As an illustration, we show those of the EC1 and EC2 clans from the Dominica dataset [13]. Only the most numerous coda types are shown: the rest of the vocal repertoires consists of more coda types with residual frequencies. C The subcoda structure can be modeled by considering rhythmic variations within codas of the same coda type. To do so, codas are represented as sequences of discrete inter-click intervals (dICIs), by discretizing absolute ICIs into discrete bins, which can then be considered akin to symbols (e.g. A, B, C, …), providing a tokenization of codas. Different instances of a single coda type can correspond to slightly different dICI sequences. The resulting dICI sequences are modeled using variable-length Markov chains, which can be represented as subcoda trees. These trees are built starting from the root node. Each node represetns a dICI and a dICI is a child of another in the tree if it’s presence in the affects the output probability of the parent. That is the probability of the parent depends on wether the child node has been observed previously in the sequence; see the Methods for a thorough explanation. These trees can be built for an individual speaker or for a group of speakers (here represented by colors), and capture the statistical properties of rhythm variations within codas and in the transitions between those. In other words, the tree captures a vocal style—how they say what they say. The vocal styles of different groups of sperm whales can be quantitatively compared by calculating a distance between their subcoda trees, which is a statistical distance between the dICI output distributions.
Sperm whales live in multi-tiered societies and have a complex vocal communication system [14]. They communicate through rhythmic patterns (codas) of short broadband sounds (clicks), which have traditionally been classified into a set of coda types. This classification is based on the total number of clicks and their rhythm and tempo extrapolated from the time interval between consecutive clicks: the inter-click interval (ICI) [15, 16] (Fig. 1A). For example, the 4-regular (4R2) type refers to a pattern of four evenly spaced clicks, whereas the 1+3 type refers to two clicks separated by a longer pause followed by two clicks in quick succession. Coda types are thus standardized rhythmic patterns, but individual vocalizations of a given coda type exhibit micro-variations around that pattern. In the case of sperm whales ICIs can be used instead of inter-onset intervals (IOIs) due to the very low variance in duration of the individual clicks [17, 18].
@@
By applying our modeling framework to acoustic data from the Atlantic and Pacific Oceans, we obtain two main results. First, we partition sperm whale populations into vocal style-defined clans, which we find to recapitulate the previously defined vocal repertoire clans. This confirms our method ’s validity as it reiterates previous studies’ results. Second, and crucially, we find that the vocal style of non-ID codas is more similar for more sympatric clans, i.e. clans whose territory overlaps more spatially. In contrast, we do not find an effect of sympatry on the similarity of vocal styles when studying only ID codas. This suggests that geographic overlap induces vocal styles to become more similar between clans, without jeopardizing each clan’s acoustic identity signals. Our results strengthen previous results on the use of ID codas as symbolic markers, while supporting cultural transmission and social learning of vocalizations among whales of different clans, as predicted by theoretical models [25].
@@
Click Sharp broadband pulses [26]
Inter-click interval (ICI) Time interval between two consecutive clicks
Discrete ICI (dICI) Discretization of an inter-click interval given a bin size.
Coda Stereotyped, short series of clicks produced by sperm whales in social situations [15]
Coda type Categorical classification of codas based on rhythm and tempo of coda inter-click intervals [19]
Coda frequency Usage frequency of each coda type
Coda sample Set of codas recorded during one event
Coda usage Set of coda types used by a group of whales
Identity codas Subset of coda types most frequently used by a sperm whale clan; canonically used to define vocal clans.
Non-identity codas Coda types produced by whales belonging to multiple vocal clans (i.e., all coda types that are not identity codas)
Vocal clan Higher-order level of social organization delineated based on social units that share a substantial part of their vocal repertoire [12, 23]
Vocal repertoire Coda usage and frequency combined. Description of what sperm whales vocalize.
Vocal style Encodes the rhythmic variations within codas. Description of how sperm whales emit codas
Vocal identity Concept comprising both vocal style and vocal repertoire
@@ Results
We model the internal structure of codas, in terms of rhythmic variations at the level of clicks, by using variable length Markov chains (VLMCs). Our analytical pipeline is illustrated in Fig. 1C. We build each VLMC in two main steps. We first convert codas, naturally represented as sequences of continuous, absolute, inter-click intervals (ICIs), to sequences of discrete ICIs (dICIs), by discretizing time into bins. In this way, each dICI represents a narrow range of possible ICI values. The bins have a fixed width (or resolution) $\delta t$ and thus implicitly correspond to the temporal resolution of our representation (see Methods for details on the optimal choice of $\delta t$). Note that although ICIs have units of time (seconds), dICIs are (unit-less) symbols (e.g. A, B, C, etc.), representing multiples of $\delta t$ (and so the smaller $\delta t$, the more the symbols). For example, the shortest ICIs will be mapped to the symbol A whereas longer ones will be mapped to symbols further down the alphabet. Hence, each coda (a sequence of ICIs) is mapped to a sequence of discrete symbols (a sequence of dICIs). The second part of the pipeline focuses on modeling the internal structure of codas in terms of dICI sequences. Essentially, we want to estimate transition probabilities from a dICI sub-sequence to the next dICI. A standard way would be to describe this using $k$-order Markov chain models, which model the next state (dICI) probability given a past sequence of $k$-states (dICIs). However, it is possible that specific dICIs depend on sequences of varying length. Despite we do not have direct evidence of unitary blocks in sperm whale communication, on can imagine this effect similarly to what happens with words (e.g., a word beginning with “re” can continue in more ways than one starting with “zy”). To account for the possibility that states (words, dICIs) might depend on sequences of varying length, while also retaining only the most compressed statistical representation of how codas are structured in terms of dICIs, we employ VLMCs.
@@
Figure 4: Clan overlap influences non-ID coda vocal styles Comparing the similarities of different VLMC models fit for each Pacific Ocean clan for both ID and non-ID coda samples. The y-axis represents the measured distance between the subcoda trees, and the x-axis shows the geographical clan overlap (as calculated in [12] ; see ”Measuring clan overlap” in our Methods for details). Each point represents a pairwise comparison between two clans. The effect of overlap on ID coda vocal style similarity is minimal and non-significant while the opposite is true for non-ID codas: overlapping clans produce non-ID codas with a more similar vocal style. The VLMC distances are also typically much greater for ID codas than for non-ID codas. Note that while these results are visually opposite to those reported in Hersh et al. [12], they support the same final conclusions (see “Identity and non-identity codas show different trends” for details).
@@
We find the exact opposite effect when considering vocal style. The similarity in vocal style for ID codas across clans does not depend on the level of clan overlap (Fig. 4a). In contrast, the similarity in vocal style for non-ID codas displays a clear and significant increase (i.e., decreasing subcoda tree-distance) as clan spatial overlap increases (Fig. 4b). In the Supplementary Materials (see Section $2.4.2$), we show that the same results hold at the single coda type level, in addition to the whole clan level, along with an analysis of the confidence intervals. These results imply that the internal structure of codas is more similar for groups that likely spend more time in the same space, which can be thought to be analogous to accents aligning in human populations that share the same territory [28, 29]. This also highlights the complementarity of vocal repertoire and style: the trends are different precisely because the two concepts describe different aspects of whale vocalization.
@@ Discussion
For ID codas, we show that the similarity between clan vocal styles is not affected by spatial overlap, while it has recently been shown that the similarity between clan vocal repertoires decreases with overlap (measured on the two clans’ joint set of ID codas) [12]. This means that spatial overlap does not affect how whales produce ID codas (in terms of their fine-scale rhythmic structure; our results) but does affect how often they produce them [12]. In contrast, for non-ID codas, we show that the similarity between vocal styles increases with spatial overlap between two clans, while no change was observed for vocal repertoires in previous work on the same dataset. In other words, increasing spatial overlap is correlated with more similar fine-scale rhythmic structure of non-ID codas produced by whales from different clans (our results), but does not affect how often non-ID codas are produced. Our study thus supports and nuances the results of Hersh et al. [12]. We provide further support for selection acting to produce unambiguous, recognizable identity signals in the ID codas. However, ID codas only account for 35% of the total vocalizations; the remaining 65% of codas have traditionally been lumped into a catch-all category (i.e., non-ID codas) and their function remains enigmatic (these numbers are an average over the Pacific clans, and go up to 93% for non-ID codas when counting number of coda types instead of number of codas emitted, see SM 1.1 for details). We could still discriminate among clans using non-ID coda vocal style; however, the increased similarity of non-ID coda vocal styles between clans with greater spatial overlap, as demonstrated here, suggests that non-ID codas are likely vocal cues and not identity signals like the ID codas. Accordingly, vocal repertoire and vocal style capture different and complementary information on sperm whale communication, and should be considered in tandem in future studies.
@@
The most parsimonious explanation for the observed similarity of non-ID coda vocal styles of clans with increasing spatial overlap is social learning across clan boundaries. This is remarkable, given that sperm whales belonging to different clans have rarely been observed physically interacting at sea [4]. However, that does not preclude the possibility that they are within acoustic range of each other [36] and that cross-cultural social learning opportunities arise. This explanation is compatible with (and bolsters) past work suggesting that ID and non-ID codas function differently in sperm whale communication, and further suggests that they experience different evolutionary pressures [25]. Whether social learning has facilitated stochastic (i.e., cultural drift) or deterministic (i.e., cultural selection) processes is more difficult to determine, and it is unclear whether the observed non-ID coda vocal style alignment has been neutral or adaptive [37, 31]. Importantly, these findings suggest that vocal learning in sperm whales may not be limited to vertical transmission from related adults to young kin, but that horizontal and/or oblique social learning from outside the natal social unit might also be occuring.
Vocal identity in sperm whales is thus consistent with both cultural selection on ID codas to maintain discrete signals for vocal recognition in sympatry, and social learning between clans leading to a vocal style more similar to that of other whales with which they are in acoustic contact more frequently. This highlights a more complex system of transmission in which clan identity is maintained through selection, while gradual change over time may occur within and across clans for vocalizations which do not function in social recognition and thus may create similar vocal styles. Indeed, the observed alignment of non-ID vocal styles may reflect a combination of automatic imitation, and controlled imitation, guided by social and ecological factors, as discussed in studies of social learning mechanisms [38, 39]. In the context of our findings, the similarity in non-identity codas across clans with geographic overlap may indicate an automatic component of social learning. Whales sharing the same habitat are exposed to overlapping acoustic environments, which could facilitate the development of similar vocal styles described here. However, the persistence of distinct identity codas within clans may suggests a more directed form of social learning, shaped by social goals such as clan recognition and cohesion.
@@
Our results can be expanded in multiple ways in future work. The first, and the simplest conceptually, would be to conduct the present analysis on a larger dataset. More codas would improve the quality of the statistical analyses and ensure that all codas are represented in realistic proportions for each clan. Moreover, longitudinal datasets might provide direct evidence to discriminate between the social learning hypothesis and competing ones (e.g. drift in vocal style). Similarly, confirmations could emerge from large scale genetic datasets addressing the issues of phylogenetic relatedness (or lack thereof) in clans that are closer in vocal style distance. Such datasets do not exist at present, but efforts towards automated and semi-automated collection techniques are underway (e.g. Project CETI [40]). Second, from a methodological perspective, we could add spectral information (in terms of acoustic frequencies) to the temporal information currently used. Although sperm whale acoustic communication seems mostly based on rhythm, spectral features of individual clicks may convey additional information. This possibility could be incorporated into our method by labeling the dICIs according to the frequency content of the associated click (or by extending the available “alphabet” for the VLMC). Third, it would be interesting to investigate in more detail the function of non-ID codas. Indeed, even though ID codas were only recently formally named for the first time, they have been the primary focus of sperm whale coda research for decades. As previously mentioned, non-ID codas are a catch-all category for anything that is not an ID coda, but that does not mean that all non-ID codas function in the same way. To start to unveil their function, we need to consider the context (behavioral, environmental, etc.) in which different non-ID codas are produced [41]. The pattern we documented may or may not apply to all non-ID codas, but it is at least strong enough that we detect the relationship with clan spatial overlap when collectively considering all non-ID codas. Finally, the methods we developed here could in principle be used to capture the structural complexity of any communication system, including those coming from other animal species, provided they can be encoded in sequences of tokens (e.g. calls, signs, etc.). This would also allow to compare the complexity of communication systems across species, possibly linking it to other ecological parameters of the environments inhabited by such species.